OpenStudy (anonymous):

if f(x)=2x^2 and g(x)=2x-3 calculate the following f(1)= g(-3)=

OpenStudy (anonymous):

f(1)=2*(1^2) = 2 g(-3)=2*(-3)-3 = -9 2-9=-7

OpenStudy (anonymous):

i see cheers thanks for that!